# SOPHY: Generating Simulation-Ready Objects with PHYsical Materials

### v1

@@ Abstract
We present SOPHY, a generative model for 3D physics-aware shape synthesis. Unlike existing 3D generative models that focus solely on static geometry or 4D models that produce physics-agnostic animations, our approach jointly synthesizes shape, texture, and material properties related to physics-grounded dynamics, making the generated objects ready for simulations and interactive, dynamic environments. To train our model, we introduce a dataset of 3D objects annotated with detailed physical material attributes, along with an annotation pipeline for efficient material annotation. Our method enables applications such as text-driven generation of interactive, physics-aware 3D objects and single-image reconstruction of physically plausible shapes. Furthermore, our experiments demonstrate that jointly modeling shape and material properties enhances the realism and fidelity of generated shapes, improving performance on generative geometry evaluation metrics. Our project page is available at: https://xjay18.github.io/SOPHY.
Figure 1: Left: We introduce a dataset of 3D objects enriched with physical material properties, e.g., Young’s modulus ($E$) and Poisson’s ratio ($\nu$), to capture their physical behavior under forces and interactions with other objects. Middle: We also introduce a physics-aware generative model of objects that can be used for 4D generation from a single image along with forces in an environment (e.g., an external force for the plant shown in blue, collision with the ground for the pillow. Right: We also show text-to-4D generation as another application.
@@ 1 Introduction
Generating high-quality 3D assets that can be incorporated in interactive virtual worlds or manufactured into functional 3D objects is a fundamental challenge in generative AI and digital content creation. Unfortunately, current generative models of 3D shapes are limited to generating only static geometry and textures. Although there is significant effort in building 4D generative models for both geometry and motion synthesis, current approaches are limited to generating fixed animations and are incapable of producing motions resulting from complex, dynamic object interactions.
@@
In this work, we introduce a diffusion-based generative model for physics-aware 3D object synthesis, capable of generating 3D assets with detailed shape, texture, and, most importantly, physical material attributes governing kinematic object deformations (e.g., elastic deformations, plastic softening) and frictional interactions. Our approach produces physics-aware, interactive 3D objects, which are helpful for downstream applications such as physically based simulation, robotic interaction, and manufacturing. There are various challenges to developing such a generative approach. First, current 3D and 4D datasets do not contain physical material attributes for objects, making it hard to train, or even fine-tune such physics-aware generative models. Second, it is unclear how to jointly model the interplay of shape and material in 3D asset generation. Clearly, material attributes are not independent of geometric shapes i.e., certain materials are associated with certain geometric structures, as also repeatedly observed in prior works of physical material prediction from shapes  [26, 28, 48, 1].
@@
Our work aims to address the above challenges. First, we introduce a semi-automatic pipeline for annotation, which combines VLM guidance with iterative expert (mechanical engineer) feedback for more efficient object annotation with material properties. Second, we introduce a diffusion-based generative model that builds upon recent advancements in latent diffusion models, yet incorporates novel insights on representing 3D objects as compact latent codes capturing shape geometry, albedo color, and material properties related to elasticity and plasticity, as well as capturing their interdependence through cross-attention blocks. We observe that modeling shape and material properties together enhances the realism and fidelity of the geometry in the generated shapes. Overall, our contributions are:
@@
  • We introduce a 3D dataset consisting of 3K objects and 15K parts annotated with physical material properties.
@@
  • We propose a novel autoencoder for representing physics-aware 3D objects into compact latent codes, along with a latent diffusion-based generative model that jointly synthesizes geometry, texture, and physical materials.
  • We demonstrate applications of our generative framework on the synthesis of simulation-ready objects conditioned on text prompts or images.
@@ 2 Related Work
Over the recent years, we have witnessed an explosion of 3D generative models capable of generating detailed geometry and texture. The advances have been driven by expressive neural representations, such as NeRFs  [35, 41], Gaussian splats [20], signed distance fields [8, 39], occupancy fields [34], and numerous other representations for capturing shape or/and texture [36, 46, 14, 27, 18, 62, 59, 40, 11, 13] to name a few. We also refers readers to the recent surveys of 3D generative model [25, 56]. Unfortunately, the vast majority of existing 3D generative models neglect the physical material properties of the generated 3D objects, limiting their applicability to interactive and simulation-based environments. To address this gap, recent research has begun incorporating physical priors and constraints into the generative pipeline. For example, DiffuseBot [55] evaluates generated 3D robot models based on their simulation performance and refines the sampling distribution to favor more successful designs. Atlas3D [7] and Phys-Comp [12] enforce static equilibrium constraints during shape optimization, ensuring that generated objects remain stable under gravity. Despite these advances, all these prior works assume uniform or limited material properties, restricting their ability to generate diverse and physically intricate 3D assets. In contrast, our approach jointly optimizes 3D shape, texture, and material properties with a network that captures their natural dependencies.
@@
The advances in 3D generative models have also been led by the development of influential 3D datasets such as ShapeNet [5] and Objaverse [10]. More recent datasets, including ABO [9], Matsynth [52], and BlenderVault [31], incorporate surface texture information to train models for 3D generation and more plausible, photorealistic rendering. Despite this progress, a critical gap remains in the availability of detailed physical material information essential for accurately modeling material properties for physics-based simulations. While datasets such as ShapeNet-Mat [28], 3DCoMPaT [26], and its extensions [48, 1] attempt to address this limitation by providing part-level material labels, their annotations are coarse and unsuitable for direct use in physical simulators. To bridge this gap, we introduce a new dataset consisting of 3K objects spanning 12 shape categories, each annotated with precise part-level physical material properties.
@@
4D generation aims to create dynamic 3D content that aligns with input conditions such as images, text, or videos. Existing approaches primarily follow a data-driven pipeline, leveraging off-the-shelf image or video diffusion models to generate dynamic scenes. For example, DreamGaussian4D [44] adopts Stable Video Diffusion [3] to animate 3D Gaussian splats (3DGS) [20] reconstructed from a single image. STAG4D [60] initializes multi-view images using image-to-image diffusion [47] anchored on input video frames from a text-to-video module. However, these methods rely on pre-trained generative models that lack physical understanding, often resulting in physically unrealistic motions [2]. To address this, PhysGaussian [58] and PhysDreamer [63] integrate the Material Point Method (MPM) [17] with 3DGS to produce physics-aware dynamics. Building on these efforts, several concurrent studies have emerged recently, focusing on image-conditioned physical dynamics generation. Phy124 [29] obtains 3DGS from a single image, feeds them into an MPM simulator, and then renders the dynamic content. In PhysMotion [50], the visual quality is further enhanced with a diffusion-based video refinement module. However, these approaches often rely on manually assigned or predefined homogeneous material properties, failing to capture the diverse and heterogeneous nature of real-world materials. They also follow a two-stage pipeline, first generating 3D shapes and then assigning material properties, which can lead to inconsistencies between shape and materials. In contrast, our method introduces an end-to-end approach for simulation-ready object generation. By jointly modeling shape geometries and material properties, our method ensures greater physical realism and geometric coherence.
@@ 3 Material Annotation
We first describe the procedure for creating a dataset of 3D objects annotated with detailed physical material properties. Specifically, we annotate shapes with the material parameters used in the Material Point Method (MPM) [49, 17, 15], a popular simulator known for its effectiveness in handling complex simulations of behaviors and inertia effects inherent in solids  [58, 24, 4, 63], including but not limited to elastic and plastic deformation of solids and frictional interactions. The material parameters include: (a) Young’s modulus, (b) Poisson’s ratio, (c) yield stress, (d) friction angle, and (e) material behavior type (also known as material model), including pure elastic deformation [57], plastic deformation with softening (damage) [53], plastic deformation without softening [54], and granular deformation [22].
Unfortunately, no datasets of 3D objects exist with such detailed material properties available. The most related dataset is 3DCoMPaT [26] along with its subsequent versions (3DCoMPaT++ [48], 3DCoMPaT200 [1]), which provide a rather coarse labeling of materials for 3D shape parts, such as “plastic”, “fabric”, “metal”, “wood”, and so on. While these classifications provide some insight into the general physical characteristics of objects, they are often too broad for accurate simulation in physics engines. For example, a “plastic” part made out of rigid PVC is much stiffer (i.e., has a much higher Young’s modulus) than a flexible plastic material (e.g., LDPE). Unfortunately, such properties can be provided only by domain experts or manufacturers, and are difficult to interpret or estimate for annotators without a strong physics background and experience on commonly used material types in objects. Hiring experts to label material properties for every single component of every 3D object in a large database would be extremely cumbersome. Thus, we devised a semi-automatic pipeline that combines material property annotation by VLMs, followed by iterative expert feedback and verification.
@@ 3.1 VLM-guided Material Property Proposals
VLMs have recently become prominent in physical reasoning [65, 6, 23, 64, 30, 32] due to their extensive knowledge bases built on multi-modal data. Thus, we leveraged their zero-shot reasoning ability to give an initial estimation of material properties for input 3D shapes. We started by choosing $12$ shape categories from 3DCoMPaT200 [1], which contained a variety of material compositions and behaviors capable of elastic or plastic deformation, such as bags, pillows, and chairs. We skipped categories where objects behave rigidly (e.g., cabinets, faucets). As shown in Figure 2(top), we provided a popular VLM (ChatGPT-4o [37]) with (a) two automatically generated, rendered views of each textured object in the selected categories, (b) a textual description of the shape, including its object category, part tag, coarse material label, and color, as provided by 3DCoMPaT200, (c) a list of available fine-grained material categories e.g., for “plastic”, we include sub-types such as polypropylene, rigid or flexible PVC, and so on. We provide this fine-grained list in Section 7.1 (supplement). Our text prompt asks the VLM to identify the most plausible fine-grained material category per part. Then for each part, in a second round, we further prompted the VLM to provide the most likely material models and parameters (e.g., Young’s modulus, Poisson’s ratio, yield stress, and friction angle) based on a similar textual description, including this time the fine-grained material category as shown in Figure 2(bottom). This two-round prompt provided better material property prediction based on our expert verification.
@@ 3.2 Expert Verification and Feedback
Specifically, we created five test scenarios to simulate object dynamics: (1) Dropping an object from a certain height; (2) Throwing an object in a certain direction; (3) Tilting an object; (4) Dragging an object; (5) Applying a short-term, time-variant (e.g., wind) force. We render objects photorealistically under these scenarios using a particle-based simulator based on warp-mpm [66] that we extended it to model heterogeneous materials per object, since our dataset often contains objects whose components are made of different materials. Then we used PhysGaussian [58] to render the simulation videos. After collecting these videos, we enlisted five mechanical engineer graduates to evaluate the physical plausibility of each video and instructed them to reach a consensus on their judgment. As shown in Figure 3(2)-(3a), for simulated objects deemed as realistic by this group, we store their material properties in our dataset. For objects deemed to have unrealistic motion, we ask the group to provide feedback to the VLM, specifying which part of the object seems to move in an implausible manner and why. This allows us to re-query VLMs for material parameters, as demonstrated in Figure 3(3b.1)-(3b.5), and generate new simulation videos automatically based on the VLM’s updated parameters. This process is iterative until satisfactory simulation results are achieved.
Figure 3: Expert verification of material annotations.
Figure 4:  SOPHY’s autoencoder architecture for compressing 3D objects into compact shape, texture, and material-aware codes.
@@ 3.3 Dataset Summary
After the above verification stage, we obtained $3,004$ 3D models in $12$ object categories originating from the 3DCoMPaT200 [1] dataset. Every shape part is labeled with physical material properties and detailed material categories, resulting in a total of $15,575$ labeled parts. We split our dataset into training, validation, and test sets, following the 3DCoMPaT200 splits – the total number of samples in each split are $2462$, $180$, and $362$ respectively. More information about our dataset and samples of simulation sequences can be found in Section 8 of our supplement.
@@ 4 Generative Model
Our proposed generative model, SOPHY, is based on the idea of Stable Diffusion [45], which generates data by denoising a compressed feature space (latent space). To extract this latent space, we process our shapes along with the color and physical material parameters through a variational autoencoder (Section 4.1). Then a latent diffusion model is trained to jointly model the distribution of object geometries, colors, and material properties in this latent space (Section 4.2). The rationale behind jointly modeling these properties together is that they are strongly interdependent on each other i.e., certain shapes are strongly correlated with the use of specific materials and colors e.g., thin cantilever-shaped chair bases are associated with the use of metal and have a metallic gray color appearance. Modeling these properties independently of each other would yield unlikely or impossible materials for sampled geometries. After training our diffusion model, it can be sampled to generate new latent codes, which are then decoded to generate shapes and compatible colors and materials.
4.1 Shape, Texture & Material Autoencoder
The goal of our autoencoder is to compress input 3D objects into compact latent codes that encode shape, color (albedo texture) and physical material properties. The autoencoder architecture is shown in Figure 4. We discuss its components in the following paragraphs.

### v2

# SOPHY: Learning to Generate Simulation-Ready Objects with PHYsical Materials

@@ Abstract
We present SOPHY, a generative model for 3D physics-aware shape synthesis. Unlike existing 3D generative models that focus solely on static geometry or 4D models that produce physics-agnostic animations, our method jointly synthesizes shape, texture, and material properties related to physics-grounded dynamics, making the generated objects ready for simulations and interactive, dynamic environments. To train our model, we introduce a dataset of 3D objects annotated with detailed physical material attributes, along with an efficient pipeline for material annotation. Our method enables applications such as text-driven generation of interactive, physics-aware 3D objects and single-image reconstruction of physically plausible shapes. Furthermore, our experiments show that jointly modeling shape and material properties enhances the realism and fidelity of the generated shapes, improving performance on both generative geometry and physical plausibility. Our project page is available at: https://xjay18.github.io/SOPHY_page.
Figure 1:  SOPHY is a physics-aware generative model designed for creating 3D objects that possess plausible geometries, textures, and physical properties, making them ready for simulation. (Top) Our method can generate these assets from a single image; (middle) or from a text prompt (conditions are shown as insets in the top right corner of each generated object). (Bottom) We further showcase SOPHY’s potential for constructing a physically accurate virtual world by combining various image and text conditions to generate diverse objects.
@@ 1 Introduction
Generating high-quality 3D assets that can be incorporated in interactive virtual worlds or manufactured into functional 3D objects is a fundamental challenge in generative AI and digital content creation. Unfortunately, current generative models of 3D shapes [13, 89, 61] are limited to generating only static geometry and textures. Although there is significant effort in building 4D generative models for both geometry and motion synthesis [78, 66, 38], current methods are limited to fixed animations and are incapable of producing motions resulting from complex, dynamic object interactions.
@@
In this work, we introduce a diffusion-based generative model for physics-aware 3D object synthesis, capable of generating 3D assets with detailed shape, texture, and, most importantly, physical material attributes governing kinematic object deformations (e.g., elastic deformations, plastic softening) and frictional interactions. Our approach produces physics-aware, interactive 3D objects, which are helpful for downstream applications such as physically based simulation, robotic interaction, and manufacturing. There are various challenges to developing such a generative approach. First, current 3D and 4D datasets do not contain physical material attributes for objects, making it hard to train, or even fine-tune such physics-aware generative models. Second, it is unclear how to jointly model the interplay of shape and material in 3D asset generation. Clearly, material attributes are not independent of geometric shapes, as repeatedly observed in prior works of physical material recognition from shapes  [37, 40, 63, 1]. For example, a chair’s backrest made of thin rods or strips is usually metal, while a thicker, block-like backrest is often made of fabric or soft padding.
@@
Our work aims to address the above challenges. First, we introduce a semi-automatic annotation pipeline, combining Vision Language Model (VLM)’s guidance with iterative expert (mechanical engineer) feedback for more efficient object annotation with material properties. Second, we introduce a generative model that builds upon recent advancements in latent diffusion models, yet incorporates novel insights on representing 3D objects as compact latent codes capturing shape geometry, albedo color, and material properties related to elasticity and plasticity, as well as capturing their interdependence through cross-attention blocks. Additionally, we integrate a texture enhancement module as a post-processing step that refines the generated albedo color, resulting in a high-quality texture map for improved appearance. Overall, we observe that modeling shape and material properties together enhances the realism and fidelity of the geometry in the generated shapes. Our contributions are:
@@
  • We present a semi-automatic pipeline to obtain annotations describing object physical behavior, and a new dataset of 3K objects and 15K parts annotated with detailed physical material properties.
@@
  • We introduce an autoencoder for representing physics-aware 3D objects into compact latents, along with a diffusion-based model that jointly synthesizes geometry, texture, and physical materials.
  • We showcase applications of our framework for synthesizing simulation-ready objects based on texts or images, as well as physics-aware virtual worlds filled with objects generated by our method.
@@ 2 Related Work
Over the recent years, we have witnessed an explosion of 3D generative models capable of generating detailed geometry and texture. The advances have been driven by expressive neural representations, such as NeRFs  [47, 54], Gaussian splats [27], signed distance fields [14, 52], occupancy fields [46], and numerous other representations for capturing shape or/and texture [49, 59, 21, 39, 82, 10, 11, 25, 85, 53, 18, 20, 81, 2] to name a few. We also refer readers to the recent surveys of 3D generative model [34, 75]. Unfortunately, the vast majority of existing 3D generative models neglect the physical material properties of the generated 3D objects, limiting their applicability to interactive and simulation-based environments. To address this gap, recent research has begun incorporating physical priors and constraints into the generative pipeline. For example, DiffuseBot [74] evaluates generated 3D robot models based on their simulation performance and refines the sampling distribution to favor more successful designs. Atlas3D [12] and Phys-Comp [19] enforce static equilibrium constraints during shape optimization, ensuring that generated objects remain stable under gravity. Despite these advances, all these prior works assume uniform or limited material properties, restricting their ability to generate diverse and physically intricate 3D assets. In contrast, our approach jointly optimizes 3D shape, albedo color, and material properties with a network that captures their natural dependencies.
@@
The advances in 3D generative models have also been led by the development of influential 3D datasets such as ShapeNet [7] and Objaverse [16]. More recent datasets, including ABO [15], Matsynth [71], and BlenderVault [42], incorporate surface texture information to train models for 3D generation and more plausible, photorealistic rendering. Despite this progress, a critical gap remains in the availability of detailed physical material information essential for accurately modeling material properties for physics-based simulations. While datasets such as ShapeNet-Mat [40], 3DCoMPaT [37] and its extensions [63, 1] attempt to address this limitation by providing part-level material labels, their annotations are coarse and unsuitable for direct use in physical simulators. To bridge this gap, we introduce a new dataset consisting of 3K objects spanning 12 shape categories, each annotated with precise part-level physical material properties.
@@
4D generation aims to create dynamic 3D content that aligns with input conditions such as images, text, or videos. Existing approaches primarily follow a data-driven pipeline [32, 76, 78, 38, 33, 65], leveraging off-the-shelf image or video diffusion models to generate dynamic scenes. For example, DreamGaussian4D [57] adopts Stable Video Diffusion [5] to animate 3D Gaussian splats (3DGS) [27] reconstructed from a single image. STAG4D [84] initializes multi-view images using image-to-image diffusion [60] anchored on input video frames from a text-to-video module. However, these methods rely on pre-trained generative models that lack physical understanding, often resulting in physically unrealistic motions [3]. To address this, PhysGaussian [80] and PhysDreamer [87] integrate the Material Point Method (MPM) [24] with 3DGS to produce physics-aware dynamics. Building on these efforts, several recent studies [41, 67, 9] have focused on physical dynamics generation from a single image. They typically follow a two-stage pipeline to obtain simulation-ready objects: generating 3D shapes first and then assigning them with homogeneous physical attributes by manual selection or estimation from VLMs. Despite offering an intuitive framework, they overlook the interdependence between geometries and material properties as they split shape generation and material estimation into two steps. Consequently, the assigned attributes may not accurately reflect the material heterogeneity and the tightly coupled nature of geometry observed in the real world. In contrast, our method introduces a learning-based approach for simulation-ready object generation. By jointly modeling shape geometries and material properties in an end-to-end framework, our method ensures greater physical realism and geometric coherence.
@@ 3 Material Annotation
We first describe the procedure for creating a dataset of 3D objects annotated with detailed physical material properties. Specifically, we annotate shapes with the material parameters used in the Material Point Method (MPM) [64, 24, 22], a popular simulator known for its effectiveness in handling complex simulations of behaviors and inertia effects inherent in solids  [80, 35, 6, 87], including but not limited to elastic and plastic deformation of solids and frictional interactions. The material parameters include: (a) Young’s modulus, (b) Poisson’s ratio, (c) yield stress, (d) friction angle, (e) density, and (f) material behavior type (also known as material model), including pure elastic deformation [77], plastic deformation with softening (damage) [72], plastic deformation without softening [73], and granular deformation [29].
Unfortunately, no datasets of 3D objects exist with such detailed material properties available. The most related dataset is 3DCoMPaT [37] along with its subsequent versions (3DCoMPaT++ [63], 3DCoMPaT200 [1]), which provide a rather coarse labeling of materials for 3D shape parts, such as “plastic”, “fabric”, “metal”, etc. While these categories provide some insight into the general physical characteristics of objects, they are often too broad for accurate simulation in physics engines. For example, a “plastic” part made out of rigid PVC is much stiffer (i.e., has a much higher Young’s modulus) than a flexible plastic material (e.g., LDPE). Unfortunately, such properties can be provided only by domain experts or manufacturers, and are difficult to interpret or estimate for annotators without a strong physics background and experience on commonly used material types in objects. Hiring experts to label these properties for every single component of every 3D object in a large database would be extremely cumbersome. Thus, we devised a semi-automatic pipeline that combines material property annotation by VLMs, followed by iterative expert feedback and verification.
@@ 3.1 VLM-guided Material Property Proposals
VLMs have recently become prominent in physical reasoning [90, 8, 30, 88, 41, 43] due to their extensive knowledge bases built on multi-modal data. Thus, we leveraged their zero-shot reasoning ability to give an initial estimation of material properties for input 3D shapes. We started by choosing $12$ shape categories from 3DCoMPaT200 [1], which contained a variety of material compositions and behaviors capable of elastic or plastic deformation, such as bags, pillows, and chairs. We skipped categories where objects behave rigidly (e.g., cabinets, faucets). As shown in Figure 2(top), we provided a popular VLM (ChatGPT-4o [50]) with (a) two automatically generated, rendered views of each textured object in the selected categories, (b) a textual description of the shape, including its object category, part tag, coarse material label, and color, as provided by 3DCoMPaT200, (c) a list of available fine-grained material categories e.g., for “plastic”, we include sub-types such as polypropylene, rigid or flexible PVC, and so on. We provide this fine-grained list in Section D.1. Our text prompt asks the VLM to identify the most plausible fine-grained material category per part. Then for each part, in a second round, we further prompted the VLM to provide the most likely material models and parameters (e.g., Young’s modulus, Poisson’s ratio, yield stress, and friction angle) based on a similar textual description, including this time the fine-grained material category as shown in Figure 2(bottom). This two-round prompt provided better material property prediction based on our expert verification.
Figure 3: Expert verification of material annotations.
@@ 3.2 Expert Verification and Feedback
We created five test scenarios to simulate object dynamics: (1) Dropping an object from a certain height; (2) Throwing an object in a certain direction; (3) Tilting an object; (4) Dragging an object; (5) Applying a short-term, time-variant (e.g., wind) force. We obtain object motions under these scenarios using a particle-based simulator [91] that we extended to model heterogeneous materials per particle, since our dataset often contains objects whose components are made of different materials. Then, we render object sequences with a high-performance renderer [31, 27] and obtain the simulation videos. After collecting these videos, we enlisted five mechanical engineering graduates to evaluate the physical plausibility of each video and instructed them to reach a consensus on their judgment. As shown in Figure 3(2)-(3a), for simulated objects deemed as realistic by this group, we store their material properties in our dataset. For objects deemed to have unrealistic motion, we ask the group to provide feedback to the VLM, specifying which part of the object seems to move in an implausible manner and why. This allows us to re-query VLMs for material parameters, as demonstrated in Figure 3(3b.1)-(3b.5), and generate new simulation videos automatically based on the updated parameters. This process is iterative until satisfactory simulation results are achieved.
@@ 3.3 Dataset Summary
After the above verification stage, we obtained $3,004$ 3D models in $12$ object categories originating from the 3DCoMPaT200 [1] dataset. Every shape part is labeled with physical material properties and detailed material categories, resulting in a total of $15,575$ labeled parts. We split our dataset into training, validation, and test sets, following the 3DCoMPaT200 splits – the total number of samples in each split is $2462$, $180$, and $362$, respectively. We provide more information about our dataset, including data statistics and sample simulation sequences in Appendix C.
@@ 4 Generative Model
Our proposed generative model, SOPHY, is based on Stable Diffusion [58], which generates data by denoising a compressed feature space (latent space). To extract this latent space, we process our shapes along with the color and physical material parameters through a variational autoencoder (Section 4.1). Then, a latent diffusion model is trained to jointly model the distribution of object geometries, colors, and material properties in this latent space (Section 4.2). The rationale behind jointly modeling these properties together is that they are strongly interdependent on each other i.e., certain shapes are strongly correlated with the use of specific materials and colors e.g., thin cantilever-shaped chair bases are associated with the use of metal and have a metallic gray color appearance. Modeling these properties independently of each other would yield unlikely or impossible materials for sampled geometries. After training our diffusion model, it can be sampled to generate new latent codes, which are then decoded to obtain shapes, compatible colors and physical material properties.
Figure 4: Our pipeline for compressing objects to compact shape, texture, and material codes.
4.1 Shape, Texture and Material Autoencoder
Our autoencoder (Figure 4) compresses input 3D objects into compact latent codes that encode shape, color (albedo texture) and physical material properties. We discuss it below in more detail.